# A Dual-Function Radar-Communication System Empowered by Beyond Diagonal Reconfigurable Intelligent Surface

### v6

@@ III-A Problem Formulation
Problem $\mathcal{P}^{1}$ is a challenging non-convex problem. Particularly, the non-convexity stems from the complicated fractional SCNR expression in the objective and highly coupled optimization variables. To simplify the joint design, in the following subsection, we propose a series of transformations and an ADMM based framework to decouple problem (12) into multiple more tractable sub-problems.
@@ III-B Overview of Proposed Joint Design Framework
Based on the above proposition, the objective in problem $\mathcal{P}^{1}$ is more tractable with respect to $\{\mathbf{u}_{k}\}_{\forall k}$, $\mathbf{w}$, or $\{\bm{\phi}_{\imath}\}_{\forall\imath}$. However, it is still difficult to find the solution to $\mathcal{P}^{1}$ due to non-convex and coupled constraints (12b), (12c), and (12f). To tackle constraint (12f), we first define ${\bm{\Phi}}_{g}=[{\bm{\Phi}}_{\mathrm{T},g}^{H},{\bm{\Phi}}_{\mathrm{R},g}^{H}]^{H}$ and rewrite (12f) as ${\bm{\Phi}}_{g}^{H}{\bm{\Phi}}_{g}={\bf I}_{M}$. Then, we introduce auxiliary variables ${\bm{\Theta}}_{g}=[{\bm{\Theta}}_{\mathrm{T},g}^{H},{\bm{\Theta}}_{\mathrm{R},g}^{H}]^{H}={\bm{\Phi}}_{g}$ and decouple constraint (12f) into two separate constraints by adding the equality, which yields the following problem: [^5] By introducing new auxiliary variables $\{{\bm{\Theta}}_{g}\}$, $\mathcal{P}^{1}$ is equivalently transformed into $\mathcal{P}^{1}$, effectively decoupling the constraints and facilitating the solution.
@@ III-C Solutions to Sub-problems
The following theorem provides the solution to problem (21).
@@
Problem (21) can be transformed into a convex second-order cone programming (SOCP) problem, which can be globally solved by the interior point method (IPM).
@@
where ${{{\bf{\bar{H}}}}_{u,l}}={e^{\jmath\angle\left({{{\bf{s}}_{u}}\left[l\right]}\right)}}{\bf{Gw}}\left[l\right]{\bf{h}}_{u}^{H}$. The following theorem provides the solution to problem (22).
@@
Problem (22) can be transformed into a convex SOCP, which can be solved by IPM.
@@
Now, the remaining challenge of solving problem ${\mathcal{P}^{2}_{{\text{AL}},{{\bm{\Theta}}_{g}}}}$ lies in the orthogonal constraint (24b). A closed-form solution of problem ${\mathcal{P}^{2}_{{\text{AL}},{{\bm{\Theta}}_{g}}}}$ can be obtained based on the following theorem.
@@ III-F Convergence Analysis
Let $\left\{\left\{\mathbf{U}_{k}^{n}\right\},\mathbf{W}^{n},\left\{\mathbf{\Phi}_{g}^{n}\right\},\left\{{\bm{\Theta}}_{g}^{n}\right\}\right\}$ be a sequence generated by the algorithm 1 and assume that $\lim\limits_{n\to\infty}{\bm{\Lambda}}_{g}^{n+1}-{\bm{\Lambda}}_{g}^{n}=\mathbf{0},\forall g$. Then there exists a limit point $\left\{\left\{\bar{\mathbf{U}}_{k}\right\},\bar{\mathbf{W}},\left\{\bar{\mathbf{\Phi}}_{g}\right\},\left\{\bar{{\bm{\Theta}}}_{g}\right\}\right\}$, which is a stationary point and sub-optimal solution to (16).
@@ Appendix C Proof of Lemma 1
$=\frac{{({\bf{w}}^{n})^{H}{\bm{\Upsilon}}{{\bf{w}}^{n}}}}{{{\gamma^{n}}}}+2\Re\left\{\left[\begin{array}[]{l}\frac{{2{\bm{\Upsilon}}{{\bf{w}}^{n}}}}{{{\gamma_{n}}}}\\ \frac{{({\bf{w}}^{n})^{H}{\bm{\Upsilon}}{{\bf{w}}^{n}}}}{{(\gamma^{n})^{2}}}\end{array}\right]^{H}\left[\begin{array}[]{c}{\bf w}-{\bf w}^{n}\\ \gamma-\gamma^{n}\end{array}\right]\right\}$ (40e)

### v7

@@ III-A Problem Formulation
Problem $\mathcal{P}^{1}$ is a challenging non-convex problem. Particularly, the non-convexity stems from the complicated fractional SCNR expression in the objective and highly coupled optimization variables. To simplify the joint design, in the following subsection, we propose a series of transformations and an ADMM based framework to decouple problem $\mathcal{P}^{1}$ into multiple more tractable sub-problems.
@@ III-B Overview of Proposed Joint Design Framework
Based on the above proposition, the objective in problem $\mathcal{P}^{1}$ is more tractable with respect to $\{\mathbf{u}_{k}\}_{\forall k}$, $\mathbf{w}$, or $\{\bm{\phi}_{\imath}\}_{\forall\imath}$. However, it is still difficult to find the solution to $\mathcal{P}^{1}$ due to non-convex and coupled constraints (11b), (11c), and (11f). To tackle constraint (11f), we first define ${\bm{\Phi}}_{g}=[{\bm{\Phi}}_{\mathrm{T},g}^{H},{\bm{\Phi}}_{\mathrm{R},g}^{H}]^{H}$ and rewrite (11f) as ${\bm{\Phi}}_{g}^{H}{\bm{\Phi}}_{g}={\bf I}_{M}$. Then, we introduce auxiliary variables ${\bm{\Theta}}_{g}=[{\bm{\Theta}}_{\mathrm{T},g}^{H},{\bm{\Theta}}_{\mathrm{R},g}^{H}]^{H}={\bm{\Phi}}_{g}$ and decouple constraint (11f) into two separate constraints by adding the equality, which yields the following problem: [^5] By introducing new auxiliary variables $\{{\bm{\Theta}}_{g}\}$, $\mathcal{P}^{1}$ is equivalently transformed into $\mathcal{P}^{2}$, effectively decoupling the constraints and facilitating the solution.
@@ III-C Solutions to Sub-problems
The following theorem provides the solution to problem $\mathcal{P}_{\mathrm{AL},\mathbf{W}}^{2}$.
@@
Problem $\mathcal{P}_{\mathrm{AL},\mathbf{W}}^{2}$ can be transformed into a convex second-order cone programming (SOCP) problem, which can be globally solved by the interior point method (IPM).
@@
where ${{{\bf{\bar{H}}}}_{u,l}}={e^{\jmath\angle\left({{{\bf{s}}_{u}}\left[l\right]}\right)}}{\bf{Gw}}\left[l\right]{\bf{h}}_{u}^{H}$. The following theorem provides the solution to problem $\mathcal{P}_{\mathrm{AL},{\left\{{\bm{\Phi}}_{g}\right\}}}^{2}$.
@@
Problem $\mathcal{P}_{\mathrm{AL},{\left\{{\bm{\Phi}}_{g}\right\}}}^{2}$ can be transformed into a convex SOCP, which can be solved by IPM.
@@
Now, the remaining challenge of solving problem $\mathcal{P}_{\mathrm{AL},{{\bm{\Theta}}_{g}}}^{2-1}$ lies in the orthogonal constraint (23b). A closed-form solution of problem $\mathcal{P}_{\mathrm{AL},{{\bm{\Theta}}_{g}}}^{2-1}$ can be obtained based on the following theorem.
@@ III-F Convergence Analysis
Let $\left\{\left\{\mathbf{U}_{k}^{n}\right\},\mathbf{W}^{n},\left\{\mathbf{\Phi}_{g}^{n}\right\},\left\{{\bm{\Theta}}_{g}^{n}\right\}\right\}$ be a sequence generated by the algorithm 1 and assume that $\lim\limits_{n\to\infty}{\bm{\Lambda}}_{g}^{n+1}-{\bm{\Lambda}}_{g}^{n}=\mathbf{0},\forall g$. Then there exists a limit point $\left\{\left\{\bar{\mathbf{U}}_{k}\right\},\bar{\mathbf{W}},\left\{\bar{\mathbf{\Phi}}_{g}\right\},\left\{\bar{{\bm{\Theta}}}_{g}\right\}\right\}$, which is a stationary point and sub-optimal solution to $\mathcal{P}_{\mathrm{AL}}^{2}$.
@@ Appendix C Proof of Lemma 1
$=\frac{{({\bf{w}}^{n})^{H}{\bm{\Upsilon}}{{\bf{w}}^{n}}}}{{{\gamma^{n}}}}+2\Re\left\{\left[\begin{array}[]{l}\frac{{2{\bm{\Upsilon}}{{\bf{w}}^{n}}}}{{{\gamma_{n}}}}\\ \frac{{({\bf{w}}^{n})^{H}{\bm{\Upsilon}}{{\bf{w}}^{n}}}}{{(\gamma^{n})^{2}}}\end{array}\right]^{H}\left[\begin{array}[]{c}{\bf w}-{\bf w}^{n}\\ \gamma-\gamma^{n}\end{array}\right]\right\}$ (41)